You select two marbles from a cup with 3 black marbles and 2 marbles. Make an area model to analyze the probabilities.
a. What is the probability of selecting 2 black marbles?
b. What is the probability of selecting 2 blue marbles?
c. What is the probability of selecting a black marble and a blue marble?

Answer by stanbon(75887)

You select two marbles from a cup with 3 black marbles and 2 blue marbles. Make an area model to analyze the probabilities.
a. What is the probability (of selecting 2 black marbles) = 3C2/5C2=3/10
b. What is the probability (of selecting 2 blue marbles)=2C2/5C2=1/10
c. What is the probability (of selecting a black marble and a blue marble)
= 2(1/3)(1/2)=1.3
==========
Cheers,
Stan H.